OpenStudy (anonymous):

In triangle ABC, AB + 2(BC) = BC + 2(AB) and AC + 2(AB) = AB + 2(AC). Classify triangle ABC. Scalene Isosceles Equilateral Right

OpenStudy (anonymous):

AB + 2(BC) -AB -BC = BC + 2(AB) -AB - BC BC = AB AC + 2(AB) -AC - AB = AB + 2(AC) -AC - AB AB = AC AB = AC = BC by the transitive property therefore it must be Equilateral

OpenStudy (anonymous):

thank you so much, ytour amazing hun (: